# SOPHY: Generating Simulation-Ready Objects with PHYsical Materials

### v1

@@ 4.1 Shape, Texture & Material Autoencoder
We represent an object as a dense surface point cloud $\boldsymbol{P}=\{\boldsymbol{p}_{j}\}_{j=1}^{N}$, where $\boldsymbol{p}_{j}$ is a 3D point position ($N=2048$ in our implementation). The input to our autoencoder is a color- and material-augmented point cloud $\boldsymbol{X}=\{\boldsymbol{x}_{j}\}_{j=1}^{N}$, where each entry $\boldsymbol{x}_{j}$ contains the following per-point information:
  • the point’s 3D position $\boldsymbol{p}_{j}$.
  • the point’s RGB color $\boldsymbol{c}_{j}$.
@@
The above object’s representation is quite high-dimensional. Thus, we seek to compress it towards a more compact latent code. Following Stable Diffusion [45] and 3DShape2VecSet [62], we adopt a variational autoencoder (VAE) regularized with the KL-divergence to achieve this effect. We first use two fully connected (FC) layers to project each latent code $\boldsymbol{f}_{i}$ to mean and variance:
@@
Empirically, we achieved better results with an alternative decoder design. In this approach, we first decode geometry in terms of occupancy values, then decode color, and finally decode material properties, but only for query points classified as occupied. Performance was further improved when we explicitly split the latent codes $\boldsymbol{z}_{i}$ into three sub-codes: the shape code $\boldsymbol{z}_{i,s}$ for occupancy decoding, the color code $\boldsymbol{z}_{i,c}$ for texture decoding, and the material code $\boldsymbol{z}_{i,m}$ for material property decoding. Each sub-code has a shape of $M\times\frac{C_{0}}{3}$. To decode these components, we introduce three dedicated branches, each responsible for processing one of the sub-codes (Figure 4). We also apply cross-attention layers to the color and material branches, enabling them to attend on information from previously decoded latents. This models the natural dependency of color on geometry and material properties on both geometry and color (Figure 4). Our decoder design is inspired by the workflow commonly used in 3D asset creation, where artists typically start by defining the object’s shape, then apply textures, and finally assign physical attributes. Specifically, the cross-attention in our decoder is formulated as follows:
$\{\boldsymbol{f}^{(l)}_{i,s}\}$ $=\mathrm{SelfAttn}\big{(}\{\boldsymbol{f}^{(l-1)}_{i,s}\}\big{)},$ (4)
$\{\boldsymbol{f}^{(l)}_{i,c}\}$ $=\mathrm{CrossAttn}\big{(}\{\boldsymbol{f}^{(l-1)}_{i,c}\},\{\boldsymbol{f}^{(l-1)}_{i,s}\}\big{)},$ (5)
$\{\boldsymbol{f}^{(l)}_{i,m}\}$ $=\mathrm{CrossAttn}\big{(}\{\boldsymbol{f}^{(l-1)}_{i,m}\},\{\big{[}\boldsymbol{f}^{(l-1)}_{i,s},\boldsymbol{f}^{(l-1)}_{i,c}\big{]}\}\big{)}$ (6)
@@
Given a query point $\boldsymbol{q}\in\mathbb{R}^{3}$, the occupancy values are decoded by interpolating the geometry sub-codes from the last layer (layer $L$) and processing them through a fully connected (FC) block:
@@
Color and material properties are decoded similarly, each using their own dedicated FC block and interpolation over the corresponding color and material sub-codes. For color prediction, we apply a normalization operation after the FC block to ensure the output falls within the range $[0,1]$. Young’s modulus and yield stress are always positive, so the FC blocks predict their values on a logarithmic scale. The friction angle has a range of $[0,\pi/2]$, so we apply a sigmoid activation after the FC block and scale the output by $\pi/2$. The categorical material behavior type is predicted using a softmax activation after the FC block. We store the material properties only for query points predicted as occupied. For color, we only predict the values for query points sampled on the mesh surface obtained by marching cubes [33].
@@
While training our variational autoencoder, we jointly optimize a combination of loss functions involving 3D occupancies, colors, and material properties. Specifically, we minimize a weighted sum of the following losses: (a) occupancy loss ${\mathcal{L}}_{o}$ (binary cross-entropy), (b) color loss ${\mathcal{L}}_{c}$ ($\ell_{1}$ norm), (c) Young’s modulus loss ${\mathcal{L}}_{E}$ ($\ell_{1}$ norm applied to logarithmic values), (d) Poisson’s ratio loss ${\mathcal{L}}_{\nu}$ ($\ell_{1}$ norm), (e) yield stress loss ${\mathcal{L}}_{\sigma}$ ($\ell_{1}$ norm applied to logarithmic values), (f) friction angle loss ${\mathcal{L}}_{\phi}$ ($\ell_{1}$ norm), (g) material model loss ${\mathcal{L}}_{M}$ (cross-entropy), (h) a regularization loss ${\mathcal{L}}_{r}$ imposing a KL penalty towards a standard normal distribution on the latent codes, as typically used in VAEs [21]. We note that the color and material losses are computed only for training query points on or inside the object’s surface. For off-surface training points, we assign color and material properties by copying them from their nearest surface points. The combined loss function is:
${\mathcal{L}}={\mathcal{L}}_{o}+\sum_{\omega}\lambda_{\omega}{\mathcal{L}}_{\omega},\quad\omega=\{c,E,\nu,\sigma,\phi,M,r\},$ (8)
where $\lambda_{\omega}$ are weight coefficients listed in Section 9.3. To further enhance training, we leverage the pretrained 3DShape2VecSet model [62], which was trained on ShapeNet [5] (a larger dataset containing 55K shapes) using occupancy supervision alone. Specifically, we initialize all network layers shared with 3DShape2VecSet, including the cross-attention weights on point positions in our encoder, the self-attention weights on shape sub-codes, and the occupancy decoder weights, using their pretrained values. This provides a small boost compared to training our model from scratch on our smaller 3K-shape dataset.
@@ 4.2 Diffusion
${\mathcal{L}}_{\mathrm{edm}}={\mathbb{E}}_{\hat{\boldsymbol{Z}}\sim p_{data}}{\mathbb{E}}_{\boldsymbol{N}\sim{\mathcal{N}}(\boldsymbol{0},\sigma_{t}^{2}\boldsymbol{I})}\|D(\hat{\boldsymbol{Z}}+\boldsymbol{N},\sigma_{t},\boldsymbol{c})-\hat{\boldsymbol{Z}}\|_{2}^{2},$
where $D$ is the denoising network, $\hat{\boldsymbol{Z}}$ are training latent codes, $\boldsymbol{N}$ are added Gaussian noises, $\sigma_{t}$ denotes the noise level, and $\boldsymbol{c}$ is a signal for conditioning. We experimented with two input conditions to our denoiser: (a) an RGB image of a target object, where $\boldsymbol{c}$ represents here the extracted features from a pre-trained image encoder (“DINOv2-ViT-B/14” [38]), (b) a text prompt, where $\boldsymbol{c}$ represents features from a pre-trained text encoder (“CLIP-ViT-L/14” [43]). The denoising network consists of alternating self-attention layers for processing the latent representations and cross-attention layers for incorporating the signals $\boldsymbol{c}$. Details on the denoiser are provided in Section 9.1 of the supplement.
@@ 5 Experiments
We evaluate SOPHY on its autoencoder effectiveness (Section 5.1), its generative capabilities for image-conditioned generation (Section 5.2) and text-conditioned generation (Section 5.3) of simulation-ready 3D objects. All our comparisons against alternatives were performed in the same test split of our dataset, described in Section 3.3.
Metric Baseline w/o CA Fused SOPHY
@@
Table 1: Quantitative comparison in the auto-encoding setting. The upper, middle, and lower sections display the average metrics for predictions of material, color, and shape respectively.
@@ 5.1 Auto-encoding Evaluation
Competing approaches.
We stress that there is no existing generative model matching our setting of joint shape and physical material synthesis, thus, we compare here with alternative designs for our autoencoder:
(a) baseline is a model that excludes color and material properties from the generation process i.e., it generates a 3D shape, then predicts color conditioned on the shape through a decoder, and then the material through another decoder. The choice of this baseline attempts to answer the question of whether there is any benefit of incorporating the physical materials in the generation process i.e., whether it is simply better to generate the 3D shape first, then guess its most likely texture and material discriminatively. For this baseline, we use 3DShape2VecSet as the generative model, trained on the same split as our method. We use a texture decoder, which decodes its latent shape representation with a set of self-attention blocks to per-point colors. Then we use one more decoder to decode the latent shape representation to material properties, including a cross-attention block for conditioning on colors. The number of attention blocks and the number of parameters remain comparable to our model.
@@
For material and color prediction, we report the classification accuracy of the material behavior type (M.B. Acc.) and the mean absolute error (MAE) of all our material parameters. In addition, we report Chamfer distance (Sim-CD) between densely sampled points of the predicted shapes and ground-truth ones under the deformed states computed by our MPM simulator along the whole simulation trajectory for the dropping test scenario (Section 3.2). Finally, we report purely geometric measures for the rest state of the reconstructed test shape compared to the ground-truth one. We use the standard metrics of Chamfer Distance (CD), volumetric Intersection-over-Union (IoU), and F-score, as reported in prior 3D generative models [62]. We average all errors across the test cases of our dataset.
@@ Results.
Table 1 presents quantitative comparisons in the auto-encoding setting. Compared to the “baseline” model, jointly modeling shape, color, and material in a shared embedding space significantly improves material metrics, with additional, more modest gains in color and geometry evaluation metrics. Notably, the material behavior type classification is improved by more than $20\%$, while Sim-CD is reduced by a factor of $5$x. This performance gain supports our hypothesis that geometry, color, and material attributes are strongly correlated and should be jointly modeled in the generative process. Compared to not using cross-attention in the decoder (“w/o CA”) model, we see that our full model has $2$x lower Chamfer distance during simulation, and still slightly better performance in all other metrics. Finally, we observe that the “Fused” variant, has a tiny edge over our model in Poisson’s ratio and friction angle prediction. Yet, it performs significantly worse in terms of Sim-CD, and worse in all other measures. We suspect this behavior is due to an uneven network capacity allocation across occupancy, color, and material within a single fused latent code.
Figure 5: Qualitative comparison in the image-to-4D setting. We pick test scenarios described in Section 3.2 to create 4D videos. The description below the conditioned image illustrates the desired dynamics. The blue arrow shows the direction of the external force. In the dropping scenarios, the objects fall onto a ground plane (not shown here).
Metric DG4D [44] Baseline SOPHY
VideoPhy z-score $\uparrow$ -0.09 -0.06 0.16
DINO Similarity $\uparrow$ 0.73 0.76 0.77
Table 2: Comparison on image-conditioned 4D generation.
Metric STAG4D [60] Baseline SOPHY
VideoPhy z-Score $\uparrow$ -0.82 0.30 0.52
CLIP Similarity $\uparrow$ 0.13 0.12 0.13
Table 3: Comparison on text-conditioned 4D generation.
Figure 6: Qualitative comparison in the text-to-4D setting. The text prompts used as input condition are shown below.
@@ 5.2 Image-to-4D
In the absence of published methods for generating physics-grounded 4D scenes from a single image [^1] PhysMotion [50] is a concurrent work with no published code at the time of this submission – see related work for discussion of differences., we first compare with our “baseline” model discussed in the previous section. We also compare with a popular image-to-4D method, DreamGaussian4D, or in short DG4D  [44], which generates deforming 3D Gaussian splats (3DGS) over time conditioned on an input image, yet does not rely on any explicit physics-based representation.
@@
Given rendered images of the 4D scenes generated by any of the above compared methods, and the ones rendered from the ground-truth test objects from the same viewpoints, we leverage VideoPhy [2], a state-of-the-art VLM trained based on human annotations to evaluate whether the generated videos align with real-world physics. The method produces a per-scene score that is uncalibrated for different scenarios (i.e., it has different scales). Thus, following [50], we apply $z$-score normalization on the scores across all methods for each scene to calibrate them, allowing us to obtain a more meaningful average score across the test scenes. Positive $z$-scores mean that a method performs better than average, and negative means worse. Additionally, we assess the image similarity averaged across all frames and scenes using cosine similarity on DINO features from the pre-trained encoder of “DINOv2-ViT-B/14” [38]. Given that DG4D generates 3DGS, we also use 3DGS for rendering our generated shapes using the 3DGS renderer [20] and the splat fitting provided in PhysGaussian [58]. We note that the generation times are comparable for all methods (about $10$ minutes per scene, including the feedforward pass, simulation, and rendering, as measured on a single L40s GPU).

### v2

@@ 4.1 Shape, Texture and Material Autoencoder
The above representation is quite high-dimensional. We thus seek to compress it towards a more compact latent code. Following Stable Diffusion [58] and 3DShape2VecSet [85], we adopt a variational autoencoder (VAE) regularized with the KL-divergence to achieve this effect. We first use two fully connected (FC) layers to project each latent code $\boldsymbol{f}_{i}$ to mean and variance:
@@
Empirically, we achieved better results with an alternative design. We first decode geometry in terms of occupancy values, then decode color, and finally decode material properties, but only for query points classified as occupied. Performance was further improved when we explicitly split the latent codes $\boldsymbol{z}_{i}$ into three sub-codes: the shape code $\boldsymbol{z}_{i,s}$ for occupancy decoding, the color code $\boldsymbol{z}_{i,c}$ for texture decoding, and the material code $\boldsymbol{z}_{i,m}$ for material property decoding. Each sub-code has a shape of $M\times\frac{C_{0}}{3}$. To decode these components, we introduce three dedicated branches, each responsible for processing one of the sub-codes (see Figure 4). We also apply cross-attention layers to the color and material branches, enabling them to attend to information from previously decoded latents. This design models the natural dependency of color on geometry and material properties on both geometry and color, which is inspired by the workflow commonly used in 3D asset creation, where artists typically start by defining the object’s shape, then apply textures, and finally assign physical attributes. Specifically, the cross-attention in our decoder is formulated as follows:
$\begin{gathered}\{\boldsymbol{f}^{(l)}_{i,s}\}=\mathrm{SelfAttn}\big{(}\{\boldsymbol{f}^{(l-1)}_{i,s}\}\big{)},\ \{\boldsymbol{f}^{(l)}_{i,c}\}=\mathrm{CrossAttn}\big{(}\{\boldsymbol{f}^{(l-1)}_{i,c}\},\{\boldsymbol{f}^{(l-1)}_{i,s}\}\big{)},\\ \{\boldsymbol{f}^{(l)}_{i,m}\}=\mathrm{CrossAttn}\big{(}\{\boldsymbol{f}^{(l-1)}_{i,m}\},\{\big{[}\boldsymbol{f}^{(l-1)}_{i,s},\boldsymbol{f}^{(l-1)}_{i,c}\big{]}\}\big{)},\end{gathered}$ (4)
@@
Given a query point $\boldsymbol{q}\in\mathbb{R}^{3}$, the occupancy values are decoded by interpolating the shape sub-codes from the last layer (layer $L$) and processing them through an FC block:
@@
Color and material properties are decoded similarly, each using its own FC block and interpolation over the corresponding color and material sub-codes. For color prediction, we apply normalization after the FC block to ensure the output falls within the $[0,1]$ range. Young’s modulus and yield stress are always positive, so the FC blocks predict their values on a logarithmic scale. The friction angle has a range of $[0,\pi/2]$, so we apply a sigmoid activation after the FC block and scale the output by $\pi/2$. The categorical material behavior type is predicted using a softmax activation after the FC block. We store the material properties only for query points predicted as occupied. For color, we only predict the values for query points sampled on the mesh surface obtained by marching cubes [44].
@@
While training our variational autoencoder, we jointly optimize a combination of loss functions involving 3D occupancies, colors, and material properties. Specifically, we minimize a weighted sum of the following losses: (a) occupancy loss ${\mathcal{L}}_{o}$ (binary cross-entropy), (b) color loss ${\mathcal{L}}_{c}$ ($\ell_{1}$ norm), (c) Young’s modulus loss ${\mathcal{L}}_{E}$ ($\ell_{1}$ norm applied to logarithmic values), (d) Poisson’s ratio loss ${\mathcal{L}}_{\nu}$ ($\ell_{1}$ norm), (e) yield stress loss ${\mathcal{L}}_{\sigma}$ ($\ell_{1}$ norm applied to logarithmic values), (f) friction angle loss ${\mathcal{L}}_{\phi}$ ($\ell_{1}$ norm), (g) density loss ${\mathcal{L}}_{\rho}$ ($\ell_{1}$ norm), (h) material model loss ${\mathcal{L}}_{M}$ (cross-entropy), (i) a regularization loss ${\mathcal{L}}_{r}$ imposing a KL penalty towards a standard normal distribution on the latent codes, as typically used in VAEs [28]. We note that the color and material losses are computed only for training query points on or inside the object’s surface. For off-surface training points, we assign color and material properties by copying them from their nearest surface points. The combined loss function is:
${\mathcal{L}}={\mathcal{L}}_{o}+\sum_{\omega}\lambda_{\omega}{\mathcal{L}}_{\omega},\quad\omega=\{c,E,\nu,\sigma,\phi,\rho,M,r\},$ (6)
where $\lambda_{\omega}$ are weight parameters for balancing different terms. To further enhance training, we leverage the pretrained 3DShape2VecSet model [85], which was trained on ShapeNet [7] (a larger dataset containing 55K shapes) using occupancy supervision alone. Specifically, we initialize all network layers shared with 3DShape2VecSet, including the cross-attention weights on point positions in our encoder, the self-attention weights on shape sub-codes, and the occupancy decoder weights, using their pretrained values. This provides a small boost compared to training our model from scratch on our smaller 3K-shape dataset. We also note that we adopt the latent set representation of 3DShape2VecSet [85] in our autoencoder due to its simplicity and efficiency in compressing 3D objects. Our main insight is the joint modeling of shape, texture, and physical properties, and this concept is applicable beyond this specific representation.
Texture enhancement.
As shown in Figure 4, the decoded object $\hat{\boldsymbol{X}}$ exhibits a rather flat and blurry visual appearance due to insufficient texture resolution inherited by the vertex-based color representation. Thus, we refine its texture based on renderings of the input shape (in the auto-encoding setting) or conditioned signals (in the generative setting). Specifically, we first generate a UV map, unwrapping the generated object, and then bake its per-vertex color information into a 2D texture based on the UV layout. Subsequently, we adopt the texture enhancement method of Paint3D [83] to refine the 2D texture for improved visual quality. This process serves as an optional post-processing step for our autoencoder. In general, it takes around 1 minute to refine the texture of an object.
@@ 4.2 Diffusion
${\mathcal{L}}_{\mathrm{edm}}={\mathbb{E}}_{\hat{\boldsymbol{Z}}\sim p_{data}}{\mathbb{E}}_{\boldsymbol{N}\sim{\mathcal{N}}(\boldsymbol{0},\sigma_{t}^{2}\boldsymbol{I})}\|D(\hat{\boldsymbol{Z}}+\boldsymbol{N},\sigma_{t},\boldsymbol{c})-\hat{\boldsymbol{Z}}\|_{2}^{2},$ (7)
where $D$ is the denoising network, $\hat{\boldsymbol{Z}}$ are training latent codes, $\boldsymbol{N}$ are added Gaussian noises, $\sigma_{t}$ denotes the noise level, and $\boldsymbol{c}$ is a signal for conditioning. We experimented with two input conditions to our denoiser: (a) an RGB image of a target object, where $\boldsymbol{c}$ represents here the extracted features from a pre-trained image encoder (“DINOv2-ViT-B/14” [51]), (b) a text prompt, where $\boldsymbol{c}$ represents features from a pre-trained text encoder (“CLIP-ViT-L/14” [56]). The denoising network consists of alternating self-attention layers for processing the latent representations and cross-attention layers for incorporating the signals $\boldsymbol{c}$. Details on the denoiser architecture and conditional signals are provided in Section B.1 and Section B.2, respectively.
@@ 5 Experiments
We evaluate SOPHY on its autoencoder effectiveness (Section 5.1), its generative capabilities for image-conditioned generation (Section 5.2) and text-conditioned generation (Section 5.3) of simulation-ready 3D objects. All our comparisons were performed in the same test split of our dataset, described in Section 3.3.
Table 1: Quantitative comparison in the auto-encoding setting. The upper, middle, and lower sections display the average metrics for predictions of material, color, and shape, respectively.
Metric B. Dec. w/o CA Fused SOPHY
@@
Table 2: Image-to-4D generation. “B. Dec.” and “B. Perc.” denote the decoder-based and perception-based baselines, respectively.
Metric DG4D B. Dec. B. Perc. SOPHY
VideoPhy $\uparrow$ -0.36 -0.22 0.18 0.40
CLIP $\uparrow$ 0.71 0.73 0.74 0.77
Time (min) $\downarrow$ 10 5 7 5
Table 3: Text-to-4D generation. “B. Dec.” and “B. Perc.” denote the decoder-based and perception-based baselines, respectively.
Metric STAG4D B. Dec. B. Perc. SOPHY
VideoPhy $\uparrow$ -1.20 -0.16 0.59 0.76
CLIP $\uparrow$ 0.13 0.15 0.18 0.18
Time (min) $\downarrow$ 66 5 7 5
@@ 5.1 Auto-encoding Evaluation
Competing methods.
We stress that there is no existing generative model matching our setting of joint shape and physical material synthesis, thus, we compare with variant designs for our autoencoder:
(a) baseline, or in short “B. Dec.”, is a model that excludes color and material properties from the generation process i.e., it generates a 3D shape, then predicts color conditioned on the shape through a decoder, and then the material through another decoder. The choice of this baseline attempts to answer the question of whether there is any benefit of incorporating the physical materials in the generation process i.e., whether it is simply better to generate the 3D shape first, then guess its most likely texture and material discriminatively. For this baseline, we use 3DShape2VecSet as the generative model, trained on the same split as our method. We use a texture decoder, which decodes its latent shape representation with a set of self-attention blocks to per-point colors. We use one more decoder to decode the latent shape representation to material properties, including a cross-attention block for conditioning on colors. The number of total parameters remains comparable to our model.
@@
For material and color prediction, we report the classification accuracy of the material behavior type (M.B. Acc.) and the mean absolute error (MAE) of all our material parameters. In addition, we report Chamfer distance (Sim-CD) between densely sampled points of the predicted shapes and ground-truth ones under the deformed states computed by our MPM simulator along the whole simulation trajectory for the dropping test scenario (Section 3.2). Finally, we report purely geometric measures for the rest state of the reconstructed test shape compared to the ground truth. We use the standard metrics of Chamfer Distance (CD), volumetric Intersection-over-Union (IoU), and F-score, as used in prior work [85]. We average all values for each object in our test set.
Figure 5: Qualitative comparison in the image-to-4D setting. The gray texts next to the conditioned images describe the desired dynamics. Note that the objects fall onto a ground plane, which is not shown here for clarity.
@@ Results.
Section 5 presents quantitative comparisons in the auto-encoding setting. Compared to the baseline model, jointly modeling shape, color, and material in a shared embedding space significantly improves material metrics, with additional, more modest gains in color and geometry evaluation metrics. Notably, the material behavior type classification is improved by more than $20\%$, while Sim-CD is reduced by a factor of 5×5$\times$5 ×. This performance gain supports our hypothesis that geometry, color, and material attributes are strongly correlated and should be jointly modeled in the generative process. Compared to not using cross-attention in the decoder (“w/o CA”) model, we see that our full model has 2×2$\times$2 × lower Chamfer distance during simulation, and still slightly better performance in all other metrics. Finally, we observe that the “Fused” variant, has a tiny edge over our model in Poisson’s ratio and friction angle prediction. Yet, it performs significantly worse in terms of Sim-CD, and worse in all other measures. We suspect this behavior is due to an uneven network capacity allocation across occupancy, color, and material within a single fused latent code.
@@ 5.2 Image-to-4D
We first compare with a popular image-to-4D method, DreamGaussian4D (DG4D) [57], which generates deforming 3DGS over time conditioned on an input image, yet does not rely on any explicit physics-based representation. We also compare with our “baseline” discussed in the previous section. In addition, we compare with another baseline inspired by recent work [41, 9], which adopts a two-stage framework for simulation-ready object generation (see Section 2), and design a training-free baseline, which we call “B. Perc”. Specifically, this alternative baseline uses the perceptual models [45, 50] for material property estimation given an off-the-shelf 3D generation model [79]. After obtaining the material properties, we use the same simulator to create dynamics as SOPHY does. We provide more details about this alternative baseline in Section B.4.
@@
Given rendered images of the 4D scenes generated by any of the above compared methods, we leverage VideoPhy [3], a state-of-the-art VLM trained based on human annotations, to evaluate whether the generated videos align with real-world physics. The method produces a per-scene score that is uncalibrated for different scenarios. Thus, following [67], we apply $z$-score normalization to calibrate the scores across all methods for each scene and report the average values. A positive $z$-score indicates a method performs better than average, while negative means worse. Additionally, we use the CLIP [56] score to evaluate the consistency between the generated results and the input conditions. We also report the average time to create a video for an object starting from an image.